OpenStudy (anonymous):

what is the limit of x->1 of In(x)?

OpenStudy (shadowfiend):

I believe ln is continuous for x > 0, so this will be the same as asking what is ln(1). \(\ln x\) is the same as \(\log_e x\)

OpenStudy (shadowfiend):

We know that: \(\log_a b = c\) means \(a^c = b\).

OpenStudy (shadowfiend):

So, we're asking what is \(\ln 1\), meaning: \[\log_e 1 = x\] We can move this around: \[e^x = 1\] And the answer to that is x = 0, since raising any number to 0 gives you 1.

OpenStudy (anonymous):

okay. thank you

OpenStudy (shadowfiend):

Thus, \[\lim_{x \rightarrow 1} \ln x =0\]

OpenStudy (shadowfiend):

No problem!